OpenStudy (eric_d):

Sketch graph of following curve, indicating points of intersection of curve with axes y=x(x+1)(x-2) @Hero

OpenStudy (eric_d):

Sketch graph of following curve, indicating points of intersection of curve with axes

hero (hero):

Sketching the graph of a function is usually a very involved process, but let's just take it one step at a time. First, we will find the x and y intercepts. @eric_d, Do you know how to find the x-intercepts of the given equation?

OpenStudy (eric_d):

x-inter= .. when y=0

hero (hero):

So when y = 0, we'll have 0 = x(x + 1)(x - 2) right? Now what can we do from here to finish finding the x-intercepts?

OpenStudy (eric_d):

yes, so, x=0,-1 and 2

OpenStudy (eric_d):

?

hero (hero):

Very good. So on your xy plane, you'll need to plot the points (0,0), (-1,0) and (2,0) correct?

OpenStudy (eric_d):

yes, I gt it Thanks @Hero

hero (hero):

Next find the y-intercept. Any idea how to find that?

OpenStudy (eric_d):

y-inter=0

hero (hero):

Yes y = 0 since 0(0 - 1)(0 - 2) = 0

OpenStudy (eric_d):